# Acceptor and donor impurity levels in hexagonal-diamond silicon

### v1

@@ IV Formation Energy
Looking at Figure 1, one can evince, as expected, that negatively charged acceptors are more stable than neutral impurities because of their tendency to attract an electron from the host crystal to create a hole in the valence band. Interestingly, for both the neutral and charged cases, their formation energy is always lower in the 2H phase with respect to the 3C one (see Figure 1). This is even clearer if one looks at the first column of Table 4 where the difference in formation energy, $\Delta E^{2H-3C}$ for a given impurity in the 3C and in the 2H phase is reported for both the charged and the neutral cases. The more $\Delta E^{2H-3C}$ is negative, the more the impurity will prefer the 2H phase with respect to the 3C-one.
@@
Our results clearly show that, close to the 2H/3C homo-junction Vincent et al. (2014); Tang et al. (2017); Amato et al. (2019), the $\Delta E_{form}^{2H-3C}$ of charged donors is positive while for charged acceptors it is negative. To align the Fermi levels of the two phases a space charge will be generated and thus the dopants will be necessarly charged at the interface. It is hence expected that a gradient in the dopant concentration at the interface between the two phases can be build up with donors accumulating on the 3C-Si side and acceptors in 2H-Si.
@@ V Thermodynamic transition levels
A similar analysis can be performed for donors. The difference of the transition energies between the two phases is about hundreths of meV (see sixth column of Table 5 and Fig. 2) and are always negative. This means that the donor dopant levels will be always shallower in 3C-Si with respect to 2H-Si. As in the case of acceptors, even here we can recognize a relation with the ionic radius of the dopant. Indeed, $\Delta\epsilon_{D}^{2H-3C}$ decreases as a function of the radius mismatch. These findings agree with the hydrogenic model of donors that considers four covalent bonds and an unpaired one whose energy relative to the bottom of the conduction band is directly proportional to the effective mass and inversely proportional to the dielectric function. As reported in Table 1, the in-plane component of the dielectric function is lower for 2H- than for 3C-Si. This means that the transition level of donors in 2H-Si will be deeper with respect to those of 3C-Si. The discussed behavior, with the appropriate differences, was already observed in polytypes of GaAs, GaP, and InP Giorgi et al. (2020).
@@
Nevertheless a note of caution is needed when looking at the transition levels shown in Figure 1 and Figure 2. Indeed, some of these levels are resonant with the crystal host band states. This is directly related to the use of total energies obtained from semi-local functionals which cannot provide accurate ionization energies for shallow dopants Wang (2009); Zhang et al. (2013) but can only qualitatively reproduce the experimental chemical trend among different type of dopants Picus et al. (1956); Burstein et al. (1956). More advanced methods, such as quasi-particle G0W0 calculations, are known to provide accurate results but they are associated with a very high computational cost. Although in our case the absolute values of the transition energies for each phase taken separately may not be correct, the calculated semi-quantitative variation between the two phases is still valid and insightful. To further improve the comparison of the transition energy levels between the two phases, we calculated them with respect to a common reference point, i.e. the vacuum level, $\Phi$, calculated from a Si slab model, as described by Eq. (3) (see SM for more details about the procedure). Indeed, even when the theoretical description is improved with beyond-DFT methods, band edges in semiconductors undergo significant shifts but charge transition levels of defects remain practically unaffected with respect to their semi-local energy positions Alkauskas and Pasquarello (2011); Komsa et al. (2010); Alkauskas et al. (2008). Thus the calculated $\Delta\epsilon_{A}^{2H-3C}$ and $\Delta\epsilon_{D}^{2H-3C}$ presented in Table 4 and Table 5 and graphically shown in Fig. 3 are a valid reference for future experiments aiming to measure the variation in the ionization energy among the two phases. In agreement with other studies Belabbes et al. (2022); Kaewmaraya et al. (2017), we found that the 3C/2H Si junction presents a type-II alignment where the difference between hexagonal and cubic valence bands maximum (VBM), $\Delta VBM^{2H-3C}\approx 0.154$ eV and between hexagonal and cubic conduction bands minimum (CBM), $\Delta CBM^{2H-3C}\approx 0.036$ eV. Thus, the 3C-VBM lies below the 2H-VBM with a difference that is of a similar order of magnitude of the calculated $|\Delta E^{2H-3C}_{form,q}|$. In particular, when the two phases are compared, donors transition levels of both materials (referred to the vacuum level) are located almost at the same energy position due to the difference in the band offset. On the other hand, acceptors in the 3C-Si will be lower compared to 2H-Si.
@@ VI Band-offset model of doping
In order to further confirm our DFT findings and to depict a full physical picture of the dopant stability in 2H and 3C-Si, we build up a simple model following the arguments employed by Dalpian et al. Dalpian et al. (2006); Dalpian and Wei (2004) to explain the stability of III-V polytypes and recently adopted by Cai et al. Cai et al. (2021) to study 2H-Ge. Ideal doping can be represented by simply adding electrons and holes to the pristine supercell of 2H- and 3C-Si taking into account total energy correction due to finite size effects Freysoldt et al. (2009). Figure 4 reports the difference in the total energy between the two phases, $\Delta E_{total}^{2H-3C}$ per atom as a function of the carrier concentration. The results show that introducing electrons tends to stabilize the 3C phase while introducing holes is a way to make the 2H phase more stable. These results are a consequence of the type II band offset between the two phases: since 2H-Si has a higher valence band maximum than 3C (see Fig. 3), it will cost less energy to create holes in the 2H-phase than in the 3C-one. On the other hand, since 3C-Si has a lower conduction band minimum than 2H-Si (see Fig. 3), it will cost less energy to add electrons in the 3C-Si than in 2H-Si. Moreover, as the 2H/3C conduction band offset is very tiny, the slope of the electron doping curve (red line of Fig. 4) is found to be small while that of the hole incorporation is larger (green line of Fig. 4). Interestingly, this fully confirms our DFT results for impurity doped supercells which show that i) acceptors are more stable in the 2H phase and have shallower transition energy levels if compared to 3C-Si, ii) donors prefer instead to occupy cubic crystal site with shallower transition energies if compared to 2H-Si.
@@ VII Conclusion
All these results assume a particular importance for future experiments. Being accurate theoretical calculations of transition energy levels in 2H-Si, these findings represent a valid theoretical reference for future measurements of the variation in the ionization energy of a given dopant among the two phases. Moreover, in the case of 2H/3C homo-junctions Vincent et al. (2014); Tang et al. (2017); Amato et al. (2019), because of the different stability of charged dopants, a gradient in dopant concentration at the interface between the two phases may be experimentally detected with donors accumulating on the 3C-Si side and acceptors in 2H-Si.
@@ Acknowledgements.
All the authors acknowledge the ANR AMPHORE project (ANR-21-CE09-0007). Part of the high-performance computing resources for this project were granted by the Institut du développement et des ressources en informatique scientifique (IDRIS) under the allocations AD010914974 and AD010915077 via GENCI (Grand Equipement National de Calcul Intensif). The authors also acknowledge the use of the CERES high performance computer cluster at the Laboratoire de Physique des Solides of the Université Paris-Saclay (Orsay, France).

### v2

@@ IV Formation Energy
Looking at Figure 1, one can see, as expected, that negatively charged acceptors are more stable than neutral impurities because of their tendency to attract an electron from the host crystal to create a hole in the valence band. Interestingly, for both the neutral and charged cases, their formation energy is always lower in the 2H phase with respect to the 3C one (see Figure 1). This is even clearer if one looks at the first column of Table 4 where the difference in formation energy, $\Delta E^{2H-3C}$ for a given impurity in the 3C and in the 2H phase is reported for both the charged and the neutral cases. The more $\Delta E^{2H-3C}$ is negative, the more the impurity will prefer the 2H phase with respect to the 3C-one.
@@
Our results clearly show that, close to the 2H/3C homo-junction [4, 5, 21], the $\Delta E_{form}^{2H-3C}$ of charged donors is positive while for charged acceptors it is negative. To align the Fermi levels of the two phases a space charge will be generated and thus the dopants will be necessarly charged at the interface. It is hence expected that a gradient in the dopant concentration at the interface between the two phases can be build up with donors accumulating on the 3C-Si side and acceptors in 2H-Si.
@@ V Thermodynamic transition levels
A similar analysis can be performed for donors. The difference of the transition energies between the two phases is about hundreths of meV (see sixth column of Table 5 and Fig. 2) and are always negative. This means that the donor dopant levels will be always shallower in 3C-Si with respect to 2H-Si. As in the case of acceptors, even here we can recognize a relation with the ionic radius of the dopant. Indeed, $\Delta\epsilon_{D}^{2H-3C}$ decreases as a function of the radius mismatch. These findings agree with the hydrogenic model of donors that considers four covalent bonds and an unpaired one whose energy relative to the bottom of the conduction band is directly proportional to the effective mass and inversely proportional to the dielectric function. As reported in Table 1, the in-plane component of the dielectric function is lower for 2H- than for 3C-Si. This means that the transition level of donors in 2H-Si will be deeper with respect to those of 3C-Si. The discussed behavior, with the appropriate differences, was already observed in polytypes of GaAs, GaP, and InP [44].
@@
Nevertheless a note of caution is needed when looking at the transition levels shown in Figure 1 and Figure 2. Indeed, some of these levels are resonant with the crystal host band states. This is directly related to the use of total energies obtained from semi-local functionals which cannot provide accurate ionization energies for shallow dopants [45, 46] but can only qualitatively reproduce the experimental chemical trend among different type of dopants, as clearly shown in the last column of Table 4 and Table 5. More advanced methods, such as quasi-particle G0W0 calculations, are known to provide accurate results but they are associated with a very high computational cost. Although in our case the absolute values of the transition energies for each phase taken separately may not be correct, the calculated semi-quantitative variation between the two phases is still valid and insightful. To further improve the comparison of the transition energy levels between the two phases, we calculated them with respect to a common reference point, i.e. the vacuum level, $\Phi$, calculated from a Si slab model [47, 48] as described by Eq. (3) (see SM for more details about the procedure [34]). Indeed, even when the theoretical description is improved with beyond-DFT methods, band edges in semiconductors undergo significant shifts but charge transition levels of defects remain practically unaffected with respect to their semi-local energy positions [41, 42, 43]. Thus the calculated $\Delta\epsilon_{A}^{2H-3C}$ and $\Delta\epsilon_{D}^{2H-3C}$ presented in Table 4 and Table 5 and graphically shown in Fig. 3 are a valid reference for future experiments aiming to measure the variation in the ionization energy among the two phases.
In agreement with other studies [15, 8, 13], we found that the 2H/3C Si junction presents a type-II alignment, as shown by the difference between hexagonal and cubic valence band maximum (VBM), $\Delta VBM^{2H-3C}\approx 0.154$ eV and between hexagonal and cubic conduction band minimum (CBM), $\Delta CBM^{2H-3C}\approx 0.036$ eV. More accurate calculations [13] with the Heyd-Scuseria-Ernzerhof hybrid functional (HSE06) [49] and the modified Becke-Johnson potential combined with local-density approximation correlation (mBJLDA) [50] functional report a valence band offset of $\Delta VBM^{2H-3C}_{mBJLDA}=0.26$ eV and $\Delta VBM^{2H-3C}_{HSE}=0.31$ eV. Calculations for the hexagonal/cubic homojunction supercell using the Local Density Approximation (LDA) with the GW correction [14] estimate a value of $\Delta VBM^{2H-3C}_{GW}=0.25$ eV, slightly larger than the one we obtained. On the other hand, in the same work [14] a value of $\Delta CBM^{2H-3C}_{GW}=0.20$ eV was reported, while Keller et al. [13] estimate conduction band offsets to be $\Delta CBM^{2H-3C}_{MBJLDA}=0.07$ eV and $\Delta CBM^{2H-3C}_{HSE}=0.04$ eV, which is closer to our GGA results. It is possible to conclude that, despite the use of different computational approaches and alignment procedures, the type-II character of the band offset is observed in various studies [14, 15, 8, 13]. The 3C-VBM lies well below the 2H-VBM, with differences on the order of hundreds of meV, similar to the magnitude of the calculated $|\Delta E^{2H-3C}_{form,q}|$. When comparing the two phases, acceptor transition levels of both materials (referred to the vacuum level) are energetically lower in 3C-Si compared to those in 2H-Si. For donors, the band offset will lead to a closer alignment of the energy levels, although their exact relative positions may slightly differ slightly depending on the computational approach used.
@@ VI Band-offset model of doping
In order to further confirm our DFT findings and to depict a full physical picture of the dopant stability in 2H and 3C-Si, we build up a simple model following the arguments employed by Dalpian et al. [17, 18] to explain the stability of III-V polytypes and recently adopted by Cai et al. [20] to study 2H-Ge. Ideal doping can be represented by simply adding electrons and holes to the pristine supercell of 2H- and 3C-Si taking into account total energy correction due to finite size effects [30]. Figure 4 reports the difference in the total energy between the two phases, $\Delta E_{total}^{2H-3C}$ per atom as a function of the carrier concentration. The results show that introducing electrons tends to stabilize the 3C phase while introducing holes is a way to make the 2H phase more stable. These results are a consequence of the type II band offset between the two phases: since 2H-Si has a higher valence band maximum than 3C (see Fig. 3), it will cost less energy to create holes in the 2H-phase than in the 3C-one. On the other hand, since 3C-Si has a lower conduction band minimum than 2H-Si (see Fig. 3), it will cost less energy to add electrons in the 3C-Si than in 2H-Si. Moreover, as the 2H/3C conduction band offset is very tiny, the slope of the electron doping curve (red line of Fig. 4) is found to be small while that of the hole incorporation is larger (green line of Fig. 4). Interestingly, this fully confirms our DFT results for impurity doped supercells which show that i) acceptors are more stable in the 2H phase and have shallower transition energy levels if compared to 3C-Si, ii) donors prefer instead to occupy cubic crystal site with shallower transition energies if compared to 2H-Si.
@@ VII Conclusion
All these results assume a particular importance for more accurate calculations beyond the DFT framework and for future experiments. On the one hand, these results could be further assessed by hybrid-DFT or GW based calculations that could confirm the difference in transition level positions among 2H and 3C-Si by providing numerical values to be directly compared with experiments. On the other hand, being a theoretical estimation of transition energy levels in 2H-Si, these findings represent a valid theoretical reference for future measurements of the variation in the ionization energy of a given dopant among the two phases. Moreover, in the case of 2H/3C homo-junctions [4, 5, 21], because of the different stability of charged dopants, a gradient in dopant concentration at the interface between the two phases may be experimentally detected with donors accumulating on the 3C-Si side and acceptors in 2H-Si.
@@ Acknowledgements.
All the authors acknowledge the ANR AMPHORE project (ANR-21-CE09-0007) and the ANR TULIP (ANR-24-CE09-5076). Part of the high-performance computing resources for this project were granted by the Institut du développement et des ressources en informatique scientifique (IDRIS) under the allocations AD010914974 and AD010915077 via GENCI (Grand Equipement National de Calcul Intensif). The authors also acknowledge the use of the CERES high performance computer cluster at the Laboratoire de Physique des Solides of the Université Paris-Saclay (Orsay, France).